Seven times a number added to -6 is equal to twice the sum of a number and 10
Answer by TimothyLamb(4379)

7x - 6 = 2(x + 10)
7x - 6 = 2x + 20
5x = 26
x = 26/5
x = 5.2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php